Please subscribe to the ACO-status mailing list for future ACO operational advisories and alerts.

The BSP2 dislodged from the bottom on 4/11/18, 00:53 UTC, moved to 25-30 mab and is floating with the current. CTD3, BPR pressure and fluorometer data may be affected.

New ACO instruments were installed during the ACO-3 cruise on 2014/11/01. Click here for ACO-3 videos.

ALOHA Cabled Observatory
Power Supply Engineering Data
Voltage of the Stacks
Voltage (DC0) :
Voltage (DC1) :
Voltage (DC2) :
Voltage (DC3) :
Voltage (DC4) :
Voltage (DC5) :
Voltage (DC6) :
Voltage (DC7) :
105.3
0
0
102
0
0
102.2
0
V
V
V
V
V
V
V
V
Date: May 20, 2018    19:12:48 UTC
A, B bus Currents
Current A bus :
Current B bus :
Current A Regulator :
Current B Regulator :
Voltage Sensor :
Voltage Seawater return :
Line Current :
Voltage Reference :
0.915
2.062
2.8
3.5
0.382
-2.276
1.639
2.473
A
A
A
A
V
V
A
V
Date: May 20, 2018    19:12:48 UTC

 

A, B bus Voltages
Voltage A bus :
Voltage B bus :
Voltage Reference :
48.756
49.171
2.471
V
V
V
Date: May 20, 2018    19:12:48 UTC

 

ACDC Converter Temperatures
Temperature (DC0) :
Temperature (DC1) :
Temperature (DC2) :
Temperature (DC3) :
Temperature (DC4) :
Temperature (DC5) :
Temperature (DC6) :
Temperature (DC7) :
49.4
23.6
22.9
47.7
19.7
23.4
49
22
°C
°C
°C
°C
°C
°C
°C
°C
Date: May 20, 2018    19:12:48 UTC